a computer monitor has a height of 9 inches with a diagonal measurement of 15 inches. how many inches wide is the rectangular screen?

Answer by stanbon(75887)

a computer monitor has a height of 9 inches with a diagonal measurement of 15 inches. how many inches wide is the rectangular screen?
---
Draw the picture.
Draw a diagonal.
You have a rectangle with height = 9, hypotenuse = 15
and base = "x".
----------------
Use Pythagoras to solve for "x":
x^2 + 9^2 = 15^2
x^2 = 225-81
x^2 = 144
x = 12 inches (width of the screen)
==============
Cheers,
Stan H.
==============